Calculate:
1. The unemployment rate?
2. Labour force participation rate
3. Absorption rate give the following :
Labour force 17948000
Employment 13422000
Working age population 32786000
1
Expert's answer
2019-08-12T15:40:54-0400

Let unemployment rate=A

Labour force participation rate=B

Labour force absorption rate=C

Labour force=a=17948000

employment=b=13422000

Working age population=c=32786000

Then,

А = (a-b)/a *100

A = (17948000-13422000)/17948000 * 100% = 25.21%

B = a/c*100, B = 17948000/32786000 * 100% = 54.7%

C = b/c *100%

C = 13422000/32786000 * 100% = 40.9%